# On the mixed monotonicity of polynomial functions

### v1

@@ Abstract
In this paper, it is shown that every univariate polynomial function is mixed monotone globally with a polynomial decomposition function. The decomposition functions can be constructed from the Gram matrix representation of polynomial functions. The tightness of polynomial decomposition functions is discussed. An example is provided to show that polynomial decomposition functions, in addition to being global decomposition functions, can be much tighter than local decomposition functions constructed using local Jacobian bounds. Furthermore, an example is provided to demonstrate the application to reachable set over-approximation.
I Introduction
@@
Another important application of mixed monotonicity is in interval observer synthesis. Interval observers are set-based state estimators which provide a bounded interval to which the systems state is guaranteed to belong to. Interval observers are designed to ensure monotonicity of the upper and lower bounds of the interval (i.e. the upper bound always remains an upper bound and the lower bound always remains a lower bound) in addition to stability of the bounds. Mixed monotonicity is exploited in interval observer synthesis in a similar way to how it is used in reachable set prediction [3, 4, 5, 6, 7]. In these references, it is typically assumed that the decomposition function is computed from the Jacobian bounds.
@@
Polynomial functions occur frequently in control theory. In this paper, the Gram matrix representation of polynomials, where a polynomial is expressed in a quadratic form, is exploited to compute mixed monotone decompositions of polynomial functions. Using the Gram matrix representation naturally leads to semidefinite programming problems [8, 9]. In constructing a mixed monotone decomposition, the semidefinite program is related to the tightness of the decomposition function.
The contribution of this paper is that it is shown that every univariate polynomial function is mixed monotone globally with a polynomial decomposition function. Moreover, a procedure to construct the polynomial decomposition function is presented. Examples are provided which show the tightness of the derived decomposition functions and show an application to reachable set over-approximation.
II Preliminaries
II-A Notation
Let $\mathbb{R}$ denote the set of real numbers. A polynomial $p:\mathbb{R}\to\mathbb{R}$ of degree $d$ is a function that is a summation of monomials up to $d$:
$p(x)=\sum_{i=0}^{d}a_{i}x^{i}.$
[…]
The transpose of a vector $x\in\mathbb{R}^{n}$ is denoted by $x^{\top}$. The set of symmetric matrices of dimension $n\times n$ is denoted by $\mathbb{S}^{n}$. A matrix $M\in\mathbb{S}^{n}$ is positive (resp. negative) semidefinite if $x^{\top}Mx\geq 0$ (resp. $\leq 0$) for all $x\in\mathbb{R}^{n}$. The notation $M\succeq 0$ (resp. $\preceq 0$) denotes that $M$ is positive (resp. negative) semidefinite. For a vector $x\in\mathbb{R}^{n}$, $\|x\|$ denotes the Euclidean norm of $x$.
@@ II-A Notation
II-B Gram matrix representation
A polynomial $p$ of degree $d$ can be represented in a quadratic form as follows [8, 9]:
@@ II-B Gram matrix representation
and $L(\alpha)$ is a linear parameterisation of the set
=$\mathcal{L}=\left\{L\in\mathbb{S}^{\sigma}:{x^{\{\sigma\}}}^{\top}L{x^{\{\sigma\}}}=0,\forall x\in\mathbb{R}\right\}.$ (1)
@@ II-C Linear algebra
Lemma 1
Let $A\in\mathbb{S}^{n}$. There exist matrices $U,V\succeq 0$ such that $A=U-V$.
Proof:
@@
Remark 1
Note that the decomposition of a symmetric matrix into a difference of two a positive semidefinite matrices is not unique. This can easily be seen from the fact that if $A=U-V$ where $U,V\succeq 0$ , then $A=\tilde{U}-\tilde{V}$ where $\tilde{U}=U+R\succeq 0$ and $\tilde{V}=V+R\succeq 0$ for any $R\succeq 0$ of appropriate dimension.
III Background on Mixed Monotonicity
@@
The decomposition function for a mixed monotone function on a given set is in general not unique. In the aforementioned applications of mixed monotonicity it is desirable to find a decomposition function that is tight in the sense that the interval $\left[g(\underline{x},\overline{x}),g(\overline{x},\underline{x})\right]$ is the smallest interval that contains the set $\left\{f(x):\underline{x}\leq x\leq\overline{x}\right\}$. This leads to tighter over-approximations of reachable sets or interval estimates in interval observers. In [11, Theorem 1], it is shown that a tight decomposition function can be implicitly constructed from an optimization problem. For example, if $f$ is a univariate scalar function with well defined extrema, then, according to [11, Theorem 1], the following is a tight mixed monotone decomposition function:
@@
Suppose $f:\mathcal{X}\to\mathcal{T}$ is differentiable everywhere in $\mathcal{X}$ where $\mathcal{X}\subseteq\mathbb{R}$ and $\mathcal{T}\subseteq\mathbb{R}$ and $\|f^{\prime}(x)\|\leq L$ for all $x\in\mathcal{X}$. Then $g(x,y)=f(x)+L(x-y)$ is a mixed monotone decomposition function for $f$ on $\mathcal{X}$.
Proof:
This is a particular case of [10, Theorem 2]. ∎
In the sequel, the Gram matrix representation of polynomial functions and their derivatives will be exploited to construct evaluable decomposition functions for polynomial functions.
IV Mixed Monotonicity of Polynomial Functions
@@
Theorem 1
Let $p(x):\mathbb{R}\to\mathbb{R}$ be a polynomial with a derivative that has a Gram matrix representation as shown in (4). There exist polynomials $q(x):\mathbb{R}\to\mathbb{R}$ and $r(y):\mathbb{R}\to\mathbb{R}$ defined by
[…]
@@
[…]
@@
    The partial derivative $\frac{\partial g}{\partial y}(x,y)=-{y^{\{\sigma\}}}^{\top}\mathcal{V}{y^{\{\sigma\}}}$. Since $\mathcal{V}\succeq 0$, $\frac{\partial g}{\partial x}(x,y)\leq 0$ for all $(x,y)\in\mathbb{R}\times\mathbb{R}$. Therefore, $g$ is monotonically decreasing in its second argument and the third property follows.
@@
IV-A Polynomial decomposition function tightness
@@
By the mean value theorem [12, Theorem 5.10],
@@
Later on, in §V-B, it will be shown via example that different decomposition functions derived from different objective functions can have a noticeable effect on tightness of the decomposition function. From the authors experience, using the objective function (9) tends to yield tighter decomposition functions. It could be the case that other objective functions yield tighter decomposition functions for different polynomials. Therefore, in practice, one must experiment with different objective functions and check tightness in the regimes of interest.
V Examples

### v2

@@ Abstract
In this paper, it is shown that every polynomial function is mixed monotone globally with a polynomial decomposition function. For univariate polynomials, the decomposition functions can be constructed from the Gram matrix representation of polynomial functions. The tightness of polynomial decomposition functions is discussed. Several examples are provided. An example is provided to show that polynomial decomposition functions, in addition to being global decomposition functions, can be much tighter than local decomposition functions constructed using local Jacobian bounds. Furthermore, an example is provided to demonstrate the application to reachable set over-approximation.
keywords:
Polynomial Systems, Mixed Monotonicity, Reachable Sets, Interval Observers, Semidefinite Programming
thanks: This paper was not presented at any IFAC meeting. Corresponding author A.M.Tahir.
1 Introduction
@@
Another important application of mixed monotonicity is in interval observer synthesis. Interval observers are set-based state estimators which provide a bounded interval to which the systems state is guaranteed to belong to. Interval observers are designed to ensure monotonicity of the upper and lower bounds of the interval (i.e. the upper bound always remains an upper bound and the lower bound always remains a lower bound) in addition to stability of the bounds. Mixed monotonicity is exploited in interval observer synthesis in a similar way to how it is used in reachable set prediction [6, 8, 11, 10, 5].
@@
Polynomial functions occur frequently in control theory. In this paper, the Gram matrix representation of polynomials, where a polynomial is expressed in a quadratic form, is exploited to compute mixed monotone decompositions of univariate polynomial functions. Using the Gram matrix representation naturally leads to semidefinite programming problems [7, 2]. In constructing a mixed monotone decomposition, the semidefinite program is related to the tightness of the decomposition function.
The contribution of this paper is that it is shown that every polynomial function is mixed monotone globally with a polynomial decomposition function. Moreover, a procedure to construct the polynomial decomposition function is presented. Examples are provided which show the tightness of the derived decomposition functions and show an application to reachable set over-approximation.
2 Preliminaries
2.1 Notation
The set of real numbers is denoted by $\mathbb{R}$, $\mathbb{R}^{n}$ denotes the set of $n$-dimensional real vectors, $\mathbb{R}^{n\times m}$ denotes the set of real $n\times m$ matrices, and $\mathbb{N}_{0}$ denotes the set of natural numbers including 0. The $i$th element of the canonical basis for $\mathbb{R}^{n}$ is denoted by $e_{i}$.
The transpose of a vector $x\in\mathbb{R}^{n}$ is denoted by $x^{\top}$. The set of symmetric matrices of dimension $n\times n$ is denoted by $\mathbb{S}^{n}$. A matrix $M\in\mathbb{S}^{n}$ is positive (resp. negative) semidefinite if $x^{\top}Mx\geq 0$ (resp. $\leq 0$) for all $x\in\mathbb{R}^{n}$. The notation $M\succeq 0$ (resp. $\preceq 0$) denotes that $M$ is positive (resp. negative) semidefinite. For a vector $\in\mathbb{R}^{n}, x\|x\|$ denotes the Euclidean norm of $x$.
@@ 2.1 Notation
The derivative of $f(x):\mathbb{R}\to\mathbb{R}$ is denoted by $f^{\prime}(x)$ and the indefinite integral of $f(x)$ is denoted by $\int f(x)dx$.
@@
2.2 Polynomials and the Gram matrix representation
Definition 1 (Definition 2.2 in [7]).
A polynomial $p$ in $x\in\mathbb{R}^{n}$ is a finite linear combination of monomials:
$p(x)=\sum_{\alpha}c_{\alpha}x^{\alpha}=\sum_{\alpha}c_{\alpha}x_{1}^{\alpha_{1}}\dots x_{n}^{\alpha_{n}},\;\;c_{\alpha}\in\mathbb{R},$
where the sum is over a finite number of $n$-tuples $\alpha=(\alpha,\dots,\alpha_{n})$, $\alpha_{i}\in\mathbb{N}_{0}$. A polynomial is called a univariate polynomial if $n=1$. Otherwise, it is called a multivariate polynomial.
The total degree of the monomial $x^{\alpha}$ is $\sum_{i=1}^{n}\alpha_{i}$. The total degree of a polynomial is the highest degree of its monomials.
A univariate polynomial $p$ of degree $d$ can be represented in a quadratic form [^1] Note that multivariate plynomials can also be expressed using the Gram matrix representation, but in this paper the Gram matrix representation is exploited only in the univariate polynomial case. as follows [7, 2]:
@@ 2.2 Polynomials and the Gram matrix representation
and $L(\alpha)$ is a linear parameterization of the set
$\mathcal{L}=\left\{L\in\mathbb{S}^{\sigma}:{x^{\{\sigma\}}}^{\top}L{x^{\{\sigma\}}}=0,\forall x\in\mathbb{R}\right\}.$ (1)
@@ 2.3 Linear algebra
Lemma 1.
Let $A\in\mathbb{S}^{n}$. There exist matrices $U,V\succeq 0$ such that A=UVA=U-V.
Proof.
@@
Remark 1.
Note that the decomposition of a symmetric matrix into a difference of two a positive semidefinite matrices is not unique. This can easily be seen from the fact that if A=UVA=U-V where $U,V\succeq 0$ , then $A=\tilde{U}-\tilde{V}$ where $\tilde{U}=U+R\succeq 0$ and $\tilde{V}=V+R\succeq 0$ for any $R\succeq 0$ of appropriate dimension.
3 Background on Mixed Monotonicity
@@
The decomposition function for a mixed monotone function on a given set is in general not unique. In the aforementioned applications of mixed monotonicity it is desirable to find a decomposition function that is tight in the sense that the interval $\left[g(\underline{x},\overline{x}),g(\overline{x},\underline{x})\right]$ is the smallest interval that contains the set $\left\{f(x):\underline{x}\leq x\leq\overline{x}\right\}$. This leads to tighter over-approximations of reachable sets or interval estimates in interval observers. In [13, Theorem 1], it is shown that a tight decomposition function can be implicitly constructed from an optimization problem. For example, if $f$ is a univariate scalar function with well defined extrema, then, according to [13, Theorem 1], the following is a tight mixed monotone decomposition function:
@@
Suppose $f:\mathcal{X}\to\mathcal{T}$ is differentiable everywhere in $\mathcal{X}$ where $\mathcal{X}\subseteq\mathbb{R}$ and $\mathcal{T}\subseteq\mathbb{R}$ and $\|f^{\prime}(x)\|\leq L$ for all $x\in\mathcal{X}$. Then g(x,y)=f(x)+L(xy)g(x,y)=f(x)+L(x-y) is a mixed monotone decomposition function for $f$ on $\mathcal{X}$.
Proof.
This is a particular case of [12, Theorem 2]. ∎
To close this section, some other methods for constructing decomposition functions are presented which will be useful in proving the results of this paper.
3.1 Constructing decomposition functions from other decomposition functions
It will be useful throughout this paper to construct decomposition functions by considering a function to be composed from a combination of functions which are mixed monotone and have known decomposition functions.
3.1.1 Linear combinations of mixed monotone functions
One of the most useful results is the following, which guarantees that linear combinations of mixed monotone functions are mixed monotone. A decomposition function can be constructed from the decomposition functions of the functions that are linearly combined.
Lemma 2.
Let $f_{1},\dots,f_{m}:\mathcal{X}\to\mathcal{T}$ be mixed monotone functions with decomposition functions $g_{1},\dots,g_{m}:\mathcal{X}\times\mathcal{X}\to\mathcal{T}$. Then, $f(x)=\sum_{i=1}^{m}a_{i}f_{i}(x)$ where $a_{1},\dots,a_{m}\in\mathbb{R}$ is mixed monotone over $\mathcal{X}$ and
$g(x,y)=\sum_{i=1}^{m}\max(a_{i},0)g_{i}(x,y)+\min(a_{i},0)g_{i}(y,x)$ (4)
is a decomposition function for it.
Proof.
Without loss of generality, let $a_{1},\dots,a_{m}$ be arranged such that $a_{1},\dots a_{l-1}\geq 0$ and $a_{l},\dots,a_{m}<0$ where $1\leq l\leq m$. With this arrangement, (4) is equivalent to the following:
$g(x,y)=\sum_{i=1}^{l-1}a_{i}g_{i}(x,y)+\sum_{j=l}^{m}a_{j}g_{j}(y,x).$ (5)
Now it is demonstrated that $g(x,y)$ satisfies the three properties of a mixed monotone decomposition function in Definition 2:
  1. 1.
    Using substitution, $g(x,x)=\sum_{i=1}^{m}a_{i}g_{i}(x,x)=\sum_{i=1}^{m}a_{i}f_{i}(x)=f(x)$. So, $f(x)$ is “embedded” along the diagonal of $g$.
  2. 2.
    Now consider $x_{1}\geq x_{2}$. Since $g_{i}$ is a decomposition function $g_{i}(x_{1},y)\geq g_{i}(x_{2},y)$ for all $i=1,\dots m$. Therefore, $a_{i}g_{i}(x_{1},y)\geq a_{i}g_{i}(x_{2},y)$ for all $i=1,\dots l-1$. It, thus, follows that
    $\sum_{i=1}^{l-1}a_{i}g_{i}(x_{1},y)\geq\sum_{i=1}^{l-1}a_{i}g_{i}(x_{2},y).$ (6)
    Similarly, $g_{j}(y,x_{1})\leq g_{j}(y,x_{2})$ for all $j=1,\dots m$. So, $a_{j}g_{j}(y,x_{1})\geq a_{j}g_{j}(y,x_{2})$ for $j=l,\dots m$. Hence,
    $\sum_{j=l}^{m}a_{j}g_{j}(y,x_{1})\geq\sum_{j=l}^{m}a_{j}g_{j}(y,x_{2}).$ (7)
    Combining (6) and (7), yields $g(x_{1},y)\geq g(x_{2},y)$, which implies that $g$ is increasing in its first argument.
  3. 3.
    The third properties follows from similar logic to the second property.
This concludes the proof. ∎
3.1.2 Compositions of mixed monotone functions
The following is a novel result, which is that compositions of mixed monotone functions are mixed monotone, and a decomposition function can be constructed using some compositions and argument swapping.
Lemma 3.
Let $f_{1}:\mathcal{X}_{1}\to\mathcal{T}_{1}$ and $f_{2}:\mathcal{X}_{2}\to\mathcal{T}_{2}$ where $\mathcal{T}_{1}\subseteq\mathcal{X}_{2}$ be mixed monotone functions with decomposition functions $g_{1}$ and $g_{2}$, respectively. The composition $f_{1\circ 2}=f_{1}\circ f_{2}:\mathcal{X}_{1}\to\mathcal{T}_{2}$ is mixed monotone over $\mathcal{X}_{1}$ and
$g_{1\circ 2}(x,y)=g_{1}(g_{2}(x,y),g_{2}(y,x))$ (8)
is a decomposition function for it.
Proof.
Now it is demonstrated that $g_{1\circ 2}(x,y)$ satisfies the three properties of a mixed monotone decomposition function in Definition 2:
  1. 1.
    By substitution, $g_{1\circ 2}(x,x)=g_{1}(g_{2}(x,x),g_{2}(x,x))=g_{1}(f_{2}(x),f_{2}(x))=f_{1}\circ f_{2}(x)$ for all $x\in\mathcal{X}_{1}$.
  2. 2.
    Consider $x,x_{1},x_{2},y\in\mathcal{X}_{1}$ where $x_{1}\geq x_{2}$. Since $g_{2}$ is a decomposition function $g_{2}(x_{1},y)\geq g_{2}(x_{2},y)$. Since $g_{1}$ is also decomposition function,
    $g_{1}(g_{2}(x_{1},y),g_{2}(y,x))\geq g_{1}(g_{2}(x_{2},y),g_{2}(y,x)).$ (9)
    Similarly, $g_{2}(y,x_{1}))\leq g_{2}(y,x_{2})$, so
    $g_{1\circ 2}(x_{1},y)$ $=g_{1}(g_{2}(x_{1},y),g_{2}(y,x_{1}))$
    $\geq g_{1}(g_{2}(x_{1},y),g_{2}(y,x_{2})).$ (10)
    Substituting $x_{2}$ for $x$ in (9) yields:
    $g_{1}(g_{2}(x_{1},y),g_{2}(y,x_{2}))\geq$ $g_{1}(g_{2}(x_{2},y),g_{2}(y,x_{2}))$
    $=g_{1\circ 2}(x_{2},y).$ (11)
    Combining (10) and (11) yields $g_{1\circ 2}(x_{1},y)\geq g_{1\circ 2}(x_{2},y)$. So $g_{1\circ 2}$ is increasing in its first argument.
  3. 3.
    Consider $x,y,y_{1},y_{2}\in\mathcal{X}_{1}$ where $y_{1}\geq y_{2}$, so $g_{2}(x,y_{1})\leq g_{2}(x,y_{2})$. Therefore,
    $g_{1}(g_{2}(x,y_{2}),g_{2}(y,x))\geq g_{1}(g_{2}(x,y_{1}),g_{2}(y,x)).$ (12)
    Furthermore, $g_{2}(y_{2},x)\leq g_{2}(y_{1},x)$, so
    $g_{1}(g_{2}(x,y_{1}),g_{2}(y_{2},x))\geq$ $g_{1}(g_{2}(x,y_{1}),g_{2}(y_{1},x))$
    $=g_{1\circ 2}(x,y_{1}).$ (13)
    Substituting $y_{2}$ for $y$ in (12) yields
    $g_{1\circ 2}(x,y_{2})=$ $g_{1}(g_{2}(x,y_{2}),g_{2}(y_{2},x))$
    $\geq g_{1}(g_{2}(x,y_{1}),g_{2}(y_{2},x)).$ (14)
    Combining (13) and (14) yields $g_{1\circ 2}(x,y_{2})\geq g_{1\circ 2}(x,y_{1})$, which means that $g_{1\circ 2}$ is decreasing in its second argument.
This concludes the proof. ∎
4 Mixed Monotonicity of Univariate Polynomials
@@
Theorem 1.
Let $p(x):\mathbb{R}\to\mathbb{R}$ be a univariate polynomial and let its derivative be denoted using the Gram matrix representation shown in (15). There exist polynomials $q(x):\mathbb{R}\to\mathbb{R}$ and $r(y):\mathbb{R}\to\mathbb{R}$ defined by
$q(x)$ $=\int{x^{\{\sigma\}}}^{\top}\mathcal{U}{x^{\{\sigma\}}}dx+p(0),$ (16a)
$r(y)$ $=\int{y^{\{\sigma\}}}^{\top}\mathcal{V}{y^{\{\sigma\}}}dy,$ (16b)
@@
    By the definition of the (anti)derivative, $p(x)=\int p^{\prime}(x)dx+p(0)$. It then follows from (15)-(17a) and (18) that
    $p(x)$ $=\int\left({x^{\{\sigma\}}}^{\top}\left(G+L(\alpha)\right){x^{\{\sigma\}}}\right)dx+p(0)$
    $=\int\left({x^{\{\sigma\}}}^{\top}\mathcal{U}{x^{\{\sigma\}}}-{x^{\{\sigma\}}}^{\top}\mathcal{V}{x^{\{\sigma\}}}\right)dx+p(0)$
@@
    The partial derivative $\frac{\partial g}{\partial y}(x,y)=-{y^{\{\sigma\}}}^{\top}\mathcal{V}{y^{\{\sigma\}}}$. Since $\mathcal{V}\succeq 0$, $\frac{\partial g}{\partial y}(x,y)\leq 0$ for all $(x,y)\in\mathbb{R}\times\mathbb{R}$. Therefore, $g$ is monotonically decreasing in its second argument and the third property follows.
@@
4.1 Polynomial decomposition function tightness
@@
By the mean value theorem [9, Theorem 5.10],
@@
Later on, in §6.2, it will be shown via example that different decomposition functions derived from different objective functions can have a noticeable effect on tightness of the decomposition function. From the authors experience, using the objective function (20) tends to yield tighter decomposition functions. It could be the case that other objective functions yield tighter decomposition functions for different polynomials. Therefore, in practice, one must experiment with different objective functions and check tightness in the regimes of interest.
5 Mixed Monotonicity of Multivariate Polynomials
Armed with Theorem 1 and Lemmas 2 and 3, it can now be proven that products of mixed monotone functions are mixed monotone. This will be useful in proving that all multivariate polynomial functions are mixed monotone and have polynomial decomposition functions.
Lemma 4.
Let $f_{1}:\mathcal{X}\to\mathcal{T}$ and $f_{2}:\mathcal{X}\to\mathcal{T}$ where $\mathcal{T}\subseteq\mathbb{R}$ be mixed monotone functions with decomposition functions $g_{1}$ and $g_{2}$, respectively. The product $f_{1\times 2}(x)=f_{1}(x)f_{2}(x)$ is a mixed monotone function over $\mathcal{X}$ and
$g_{1\times 2}(x,y)=\frac{1}{2}g_{sq}(g_{1}(x,y)+g_{2}(x,y),g_{1}(y,x)+g_{2}(y,x))$
$-\frac{1}{2}g_{sq}(-g_{1}(x,y)+g_{2}(y,x),-g_{1}(y,x)+g_{2}(x,y))$ (21)
is a decomposition function for it, where $f_{sq}:\mathbb{R}\to\mathbb{R}$ is the squaring function $f_{sq}(z)=z^{2}$ and $g_{sq}$ is any decomposition function for it.
Proof.
The first step of the proof is to express $f_{1\times 2}(x)$ in a quadratic form as follows:
$f_{1\times 2}(x)=\begin{bmatrix}f_{1}(x)\\ f_{2}(x)\end{bmatrix}^{\top}\begin{bmatrix}0&\frac{1}{2}\\ \frac{1}{2}&0\end{bmatrix}\begin{bmatrix}f_{1}(x)\\ f_{2}(x)\end{bmatrix}$.
Using an eigendecomposition, this can be further expanded to the following:
$f_{1\times 2}(x)=\frac{1}{2}\left((f_{1}(x)+f_{2}(x))^{2}-(-f_{1}(x)+f_{2}(x))^{2}\right)$
$=\frac{1}{2}\left(f_{sq}\circ(f_{1}+f_{2})(x)-f_{sq}\circ(-f_{1}+f_{2})(x)\right).$ (22)
By Theorem 1, $f_{sq}$ is mixed monotone on $\mathbb{R}$. Hence, in (22), $f_{1\times 2}$ is expressed as a linear combination of compositions of mixed monotone functions. By Lemmas 2 and 3, this means that $f_{1\times 2}$ is mixed monotone. Since $f_{sq}$ is mixed monotone, a decomposition function $g_{sq}$ exists. The decomposition (21) can then be constructed using Lemmas 2-3. ∎
Remark 2.
Lemma 4 can be used to show that products of more than two functions are mixed monotone by induction. Suppose $f_{1},\dots f_{m}$ are mixed monotone functions. The function $f(x)=\prod_{i=1}^{m}f_{i}(x)$ is mixed monotone and this follows because, by Lemma 4, $f_{1\times 2}(x)=f_{1}(x)f_{2}(x)$ is mixed monotone and has decomposition function $g_{1\times 2}$ shown in (21). Then, using Lemma 4 again, $f_{1\times 2\times 3}(x)=f_{1}(x)f_{2}(x)f_{3}(x)$ is mixed monotone and has the following decomposition function: $g_{1\times 2\times 3}(x,y)=\frac{1}{2}g_{sq}(g_{1\times 2}(x,y)+g_{3}(x,y),g_{1\times 2}(y,x)+g_{3}(y,x))-\frac{1}{2}g_{sq}(-g_{1\times 2}(x,y)+g_{3}(y,x),-g_{1\times 2}(y,x)+g_{3}(x,y))$. This process is then repeated m3m-3 more times to construct a decomposition function for $f$.
The following theorem shows that all multivariate monomials are mixed monotone and have polynomial decomposition functions.The idea behind the proof is to use Theorem 1 and Lemma 4 to prove that monomials of multivariate polynomials are mixed monotone. It then follows from Lemma 2 that every polynomial is mixed monotone.
Theorem 2.
Every polynomial $p:\mathbb{R}^{n}\to\mathbb{R}$ is mixed monotone on $\mathbb{R}^{n}$ and has a polynomial decomposition function.
Proof.
To prove this result, it is sufficient to prove that $x^{\alpha}$ is a mixed monotone function for any $n$-tuple $\alpha=(\alpha,\dots,\alpha_{n})$, $\alpha_{i}\in\mathbb{N}_{0}$, and has a polynomial decomposition function. Then the result follows using Lemma 2.
Let $p_{d}:\mathbb{R}\to\mathbb{R}$ be $p_{d}(z)=z^{d}$ where $d\in\mathbb{N}_{0}$. By Theorem 1, $p_{d}(z)$ is mixed monotone over $\mathbb{R}$ and has a polynomial decomposition function. Therefore, using Lemma 3, $p_{d}(e_{i}^{\top}x)=p_{d}(x_{i})$ is mixed monotone over $\mathbb{R}^{n}$ and has a polynomial decomposition function $g_{i}(x,y)$.
By Lemma 4, $f_{i\times j}(x)=p_{\alpha_{i}}(e_{i}^{\top}x)p_{\alpha_{j}}(e_{j}^{\top}x)=x_{i}^{\alpha_{i}}x_{j}^{\alpha_{j}}$ is mixed monotone over $\mathbb{R}^{n}$, where $i,j\in\{1,\dots,n\}$ and $g_{i\times j}(x,y)=\frac{1}{2}g_{sq}(g_{i}(x,y)+g_{j}(x,y),g_{i}(y,x)+g_{j}(y,x))-\frac{1}{2}g_{sq}(-g_{i}(x,y)+g_{j}(y,x),-g_{i}(y,x)+g_{j}(x,y))$ is a decomposition function for it. If $g_{sq}$ is chosen to be a polynomial decomposition function for $f_{sq}$, which is always possible due to Theorem 1, and $g_{i},g_{j}$ are also chosen to be polynomials, then $g_{i\times j}(x,y)$ is a polynomial in $(x,y)$ since it is comprised of compositions and linear combinations of polynomials.
By induction (cf. Remark 2), it follows that $x^{\alpha}=\prod_{i=1}^{n}f_{\alpha_{i}}(e_{i}^{\top}x)$ is a mixed monotone function on $\mathbb{R}^{n}$ and has a polynomial decomposition function. ∎
6 Examples